# Threshold Graphs are Globally Synchronizing

### v1

@@ 6.2. Synchronous pendant extension
In this section, we present a scenario in which a group of nodes forms closed g-twins without being structural closed twins, that is called synchronous pendant extension. Specifically, consider a set $S$ that satisfies the structural assumption that the additional neighborhood of some subset $S_{1}\subseteq S$ is attached to $S_{1}$ in a leaf-like manner. The geometric assumption, on the other hand, is that this additional neighborhood synchronizes with $S_{1}$.
To prepare for the presentation and proof of Lemma 6.6, we first establish Lemma 6.5, which highlights the following fact: If there exists a synchronized, fully connected set $S$ of nodes with a synchronized pendant attached, then each node in $S$ is aligned with the vector sum of the phasors of its remaining neighbors, namely those not in $S$ nor in the pendant. From the perspective of synchronization propagation, this can be interpreted as synchronization spreading from $S$ together with its pendant to the group of all other neighbors.
@@
$W=Q\cup S\cup P.$
under the following structural assumptions:
(i) For all $i\in Q$, we have $N(i)\subseteq S$;
(ii) The subgraph $G[S]$ is a complete graph;
(iii)For any $i\in S$, its neighbourhood consists of the set $P$, its neighbours in $Q$, and the rest of the clique $S\setminus\{i\}$.
Assume further that all nodes in $Q\cup S$ are synchronized, i.e., there exists $\boldsymbol{v}$ such that $\boldsymbol{v}_{i}=\boldsymbol{v}_{j}=\boldsymbol{v},\quad\forall i,j\in Q\cup S$. Then
$\sum_{i\in P}\boldsymbol{v}_{i}\Uparrow\boldsymbol{v}\quad\text{or}\quad\sum_{i\in P}\boldsymbol{v}_{i}=\boldsymbol{0}.$
Figure 16. Illustration of Lemma 6.5. Elements in the figure follow the same conventions as in Figure 18. Moreover, a large node represents a set of nodes. An edge from a small node $A_{i}$ to a large node $S$ indicates that $A_{i}$ is connected to all nodes in $S$. The diagram depicts a configuration where a set of synchronized “false twin-like” nodes, with differing but synchronized neighbors, implies that their shared neighbors (the large white node) sum to a vector pointing in the same direction as the synchronized nodes.
@@
$\sum_{j\in Q\cup(S\setminus\{i\})\cup P}\boldsymbol{v}_{j}=\mu_{i}\boldsymbol{v}_{i}.$
Since $Q\cup S$ synchronize, we have
$(|Q|+|S|-1)\boldsymbol{v}_{i}+\sum_{j\in P}\boldsymbol{v}_{j}=\mu_{i}\boldsymbol{v}_{i},\quad\text{where}\quad\mu_{i}\geq 0.$
@@
$\sum_{j\in P}\boldsymbol{v}_{j}=\left(\mu_{i}-\left(|Q|+|S|-1\right)\right)\boldsymbol{v}_{i},\quad\text{where}\quad\mu_{i}\geq 0.$
Now, we are left to show $\mu_{i}-\left(|Q|+|S|-1\right)<0$ is not possible. Indeed, let $\boldsymbol{v}$ be an $n$-dimensional vector where the $i$-th element is one if $i\in Q\cup S$, and zero otherwise. Then
$\boldsymbol{v}^{T}H\boldsymbol{v}$ $=\sum_{i\in Q\cup S}\sum_{j\in P}\cos(\theta_{i}-\theta_{j})$ (19)
@@
$=|S|\boldsymbol{v}_{i}\cdot\sum_{j\in P}\boldsymbol{v}_{j}$
where the second equality is due to there is no connection between $Q$ and $P$. Clearly, if $\mu_{i}-\left(|Q|+|S|-1\right)<0$, then the above form is negative, and this contradicts to the stability of $\boldsymbol{\theta}$. The proof is complete. ∎
We are now ready to present the case of a synchronous pendant, which constitutes one possible configuration leading to the formation of geometric stable closed twins. This case will serve as one of the core lemmas in our proof for threshold graphs.
Consider a set of nodes $S$ attaining stable equilibrium at $\boldsymbol{\theta}$ which induces a clique [^4] Forming a clique guarantees the “closed” property required for closed twins.. $S$ can be divided into $S_{1}$ which connects only to a set $P$, and a synchronized group $S_{2}$ which in addition connects to a synchronized pendant set [^5] Specifically, $Q$ can be understood as a generalization of a leaf attached to one node, and it is well known that the presence of a leaf does not affect the stability of the state. $Q$. We will show that although $S_{1}$ and $S_{2}$ are not structurally identical, this difference, arising from the synchronous pendant, does not interfere with synchronization, in other words, the two groups behave as geometric stable closed twins.
The proof of Lemma 6.6 is based on Lemma 6.5. If we can decompose the set $P$ there into $\{A\}\cup P$, where $A$ connects to all nodes in $S$, and all nodes in $P$, then we have the following Lemma.
@@
(ii) The subgraph $G[S_{1}\cup S_{2}]$ is a complete graph;
(iii) For any $i\in S_{1}\cup S_{2}$, its neighbourhood can be decomposed into three parts: the set $P$, its neighbours inside $Q$, and the rest of the clique $S_{1}\cup S_{2}$. Formally, if we denote $Q_{i}=N(i)\cap Q$, then
$N(i)\;=\;P\cup Q_{i}\cup\bigl((S_{1}\cup S_{2})\setminus\{i\}\bigr).$
@@
$N(i)\setminus\bigl(Q\cup(S_{1}\cup S_{2})\bigr)\;=\;P.$
Assume further that all nodes in $Q\cup S_{1}$ are synchronized, i.e.,
$\boldsymbol{v}_{i}=\boldsymbol{v}_{j},\quad\forall i,j\in Q\cup S_{1}.$
Then the set $S_{1}\cup S_{2}$ forms a group of geometric stable closed twins, therefore they synchronize.
Figure 17. Illustration of Lemma 6.6. Elements follow the same conventions as in previous figures. The gray-shaded nodes are synchronized; an edge with a gray background indicates synchronization between the corresponding nodes. Each orange arc separates a small blue node from a group of nodes it connects to, indicating that the vector of the blue node is in the same direction as the sum of the vectors of the nodes on the opposite side of the arc. The figure illustrates how the synchronized block $A\cup S_{1}$ extends to include $S_{2}$ through propagation enabled by the specific local structure.
@@
$\forall i\in S_{1},\quad\boldsymbol{v}_{i}\Uparrow\sum_{k\in S_{2}\cup P}\boldsymbol{v}_{k}\quad\text{or}\quad\sum_{k\in S_{2}\cup P}\boldsymbol{v}_{k}=\boldsymbol{0}.$
@@
$\forall i\in S_{1},\quad\boldsymbol{v}_{i}\Uparrow\sum_{k\in S_{2}\cup P}\boldsymbol{v}_{j}+\sum_{k\in S_{1}\setminus\{i\}}\boldsymbol{v}_{j}\quad\text{or}\quad\sum_{k\in S_{2}\cup P}\boldsymbol{v}_{k}+\sum_{k\in S_{1}\setminus\{i\}}\boldsymbol{v}_{k}=\boldsymbol{0}.$
@@
$\forall i\in S_{1},\quad\boldsymbol{v}_{i}\Uparrow\sum_{k\in S_{1}\cup P\cup S_{2}\setminus\{i\}}\boldsymbol{v}_{k}\quad\text{or}\quad\sum_{k\in S_{1}\cup P\cup S_{2}\setminus\{i\}}\boldsymbol{v}_{k}=\boldsymbol{0}.$ (20)
@@
$\forall i\in S_{2},\quad\boldsymbol{v}_{i}\Uparrow\sum_{k\in S_{2}\setminus\{i\}}\boldsymbol{v}_{k}+\sum_{j\in S_{1}\cup P}\boldsymbol{v}_{k}\quad\text{or}\quad\sum_{k\in S_{2}\setminus\{i\}}\boldsymbol{v}_{k}+\sum_{j\in S_{1}\cup P}\boldsymbol{v}_{k}=\boldsymbol{0}$
@@
$\forall i\in S_{2},\quad\boldsymbol{v}_{i}\Uparrow\sum_{k\in S_{1}\cup P\cup S_{2}\setminus\{i\}}\boldsymbol{v}_{k}\quad\text{or}\quad\sum_{k\in S_{1}\cup P\cup S_{2}\setminus\{i\}}\boldsymbol{v}_{k}=\boldsymbol{0}.$ (21)
Combined the facts 20 and 21, we get nodes in $S_{1}\cup S_{2}$ are geometric stable closed twins. ∎
@@ 7. Main proof: the induction argument
7.1. One more key lemma
To complete the argument, we establish one more lemma prior to the proof of the main theorem.
Lemma 7.1 (Synchronization of closed twins with common neighbour).
Let $S\subseteq V$ be a set of structural closed twins in a graph $G$, and let among their common neighbourhood, there exists a node $A$ such that its neighbourhood is exactly $S$, i.e., $N(A)=S$. Assume that all nodes in $S\cup\{A\}$ attain stable equilibrium at configuration $\boldsymbol{\theta}$, that is,
$\forall i\in S\cup\{A\},\quad\mu_{i}v_{i}=\sum_{j\in N(i)}v_{j}\quad\text{ with }\quad\mu_{i}\geq 0,$
then the nodes in $S\cup\{A\}$ synchronize at $\boldsymbol{\theta}$.
Proof.
Since $S$ forms a set of structural stable closed twins, we have $i,j\in S$, we have $\boldsymbol{v}_{i}=\boldsymbol{v}_{j}$ and let us denote this common phasor as $\boldsymbol{v}$. Moreover, since $A$ only connects to the set $S$ which is synchronizing, and it attains stable equilibrium at $\boldsymbol{\theta}$, we have
$\mu_{A}\boldsymbol{v}_{A}=\sum_{j\in N(A)}\boldsymbol{v}_{j}=\sum_{j\in S}\boldsymbol{v}_{j}=|S|\boldsymbol{v}$
Thus $\boldsymbol{v}_{A}=\boldsymbol{v}$. The proof is complete. ∎
7.2. Illustrative example
In this section, we present a simple but general enough example to illustrate how to use Lemma 6.6 and 7.1 to establish global synchronization of connected threshold graphs. Specifically, to show how synchronization starts from a local substructure and eventually propagates to the entire threshold graph. The proof for all connected threshold graphs follows the same idea, with additional work needed for full generalization.
Let us take a look at the threshold graph corresponding to the sequence
$0101$
We use capital letters to represent the nodes corresponding to each bit in the sequence, including the initial starting node.
A0B1C0d1E0f1G0h1I
The corresponding threshold graph is shown as follows.
Figure 19. Threshold graph corresponding to the sequence $0101$, with vertices labeled AA to $I$ in the order of their addition.
Concluding global synchronization step by step.
To show that this graph globally synchronizes, we prove that any second order stationary point $\boldsymbol{\theta}$ of $E_{G}(\boldsymbol{\theta})$ on the threshold graph must be synchronous. Clearly, any second order stationary point $\boldsymbol{\theta}$ is also a local minimum of $E_{G}(\boldsymbol{\theta})$ or stable equilibrium of the Kuramoto model 2. Thus, the Lemmas established in Chapter 4 applies here. That is, every node attains stable equilibrium in $\boldsymbol{\theta}$, according to Lemma 4.4.
We will use Lemmas 6.6 and 7.1 repeatedly in the following illustration.
Step 1. Start with the edge between $H$ and $I$. Synchronization between nodes $H$ and $I$ can be obtained from stability condition of $H$ soely. Indeed, since $H$ attains stable equilibrium at $\boldsymbol{\theta}$, there exists $\mu_{H}\geq 0$ such that $\mu_{H}\boldsymbol{v}_{H}=\sum_{j\in N(H)}\boldsymbol{v}_{j}$ and we know that $\sum_{j\in N(H)}\boldsymbol{v}_{j}=\boldsymbol{v}_{I}$, we obtain $\boldsymbol{v}_{H}=\boldsymbol{v}_{I}$. One could also use Lemma to argue $\boldsymbol{v}_{H}=\boldsymbol{v}_{I}$, as this is the simplest case of the lemma.
Step 2. Observe that the pair of node $G$ and $I$ forms geometric stable closed twins at $\boldsymbol{\theta}$. Indeed, the only difference between $N[G]$ and $N[I]$ is that $I$ additionally has a synchronizing pendant $H$. Thus $\boldsymbol{v}_{G}=\boldsymbol{v}_{I}$. Now nodes $G,I,H$ are synchronized.
@@ 7.2. Illustrative example
Step 3. Note that node $F$ connects to $G$ and $I$ which are synchronized as established in Step 2, we have $\boldsymbol{v}_{F}=\boldsymbol{v}_{G}=\boldsymbol{v}_{I}$, according to Lemma.
Step 4. Observe that nodes $E,G,I$ induce a fully connected subgraph, and nodes $G$ and $I$ have additional synchronizing pendant, therefore, according to Lemma 6.6, three of them form geometric stable closed twins. Thus, $\boldsymbol{v}_{E}=\boldsymbol{v}_{G}=\boldsymbol{v}_{I}$.
@@
Step 4. Observe that node $D$ connects to $E,G,I$ which are synchronzed, thus $\boldsymbol{v}_{D}=\boldsymbol{v}_{E}=\boldsymbol{v}_{G}=\boldsymbol{v}_{I}$, according to Lemma.
Step 5. Observe that nodes $C,E,G,I$ induce a fully-connected subgraph, and among them, nodes $E,G,I$ admit additional synchronzing pendant $D,F,H$. Consequently $C,E,G,I$ form a group of geometric stable closed twins, and therefore they synchronize.
@@
Step 6. Now note that the node $B$ connects to a group of synchronzing nodes $C,E,G,I$. Therefore according to Lemma, node $B$ synchronzes with them as well. That is, $\boldsymbol{v}_{B}=\boldsymbol{v}_{C}=\boldsymbol{v}_{E}=\boldsymbol{v}_{G}=\boldsymbol{v}_{I}$.
Step 7. Observe that nodes $A,C,E,G,I$ induce a fully-connected subgraph, and among them, nodes $C,E,G,I$ admit additional synchronzing pendant $B,D,F,H$. Consequently $A,C,E,G,I$ form a group of geometric stable closed twins, and therefore they synchronize.
In conclusion, $\boldsymbol{\theta}$ is a synchronized state, and thus the threshold graph globally synchronizes.
7.3. Main proof
@@
Before proceeding, we introduce some notation. Any connected threshold graph can be specified by a bit sequence consisting of alternating blocks of $1$’s and 0’s. Since the graph is connected, the last bit must be $1$. We group the sequence into maximal contiguous blocks of $1$’s and 0’s. Let $U_{1},U_{2},\dots,U_{k}$ denote the blocks of $1$’s (indexed from bottom to top), and let $I_{i}$ denote the 0-block directly above $U_{i}$ (if any). Thus, the sequence takes the form either
@@
depending on whether the sequence begins with a 0-block.
Proof of Theorem 1.2.
We prove the main theorem 1.2 by induction. First, assume $\boldsymbol{\theta}$ is a second-order stationary point of $E(\boldsymbol{\theta})$. Our goal is to show $\theta_{1}=\cdots=\theta_{n}$. Let the sequence defining the threshold graph consist of maximal contiguous blocks of 1’s and 0’s, arranged as either:
$U_{1},I_{1},U_{2},I_{2},\dots,U_{k}\quad\text{or}\quad I_{0},U_{1},I_{1},U_{2},I_{2},\dots,U_{k}.$
Base case. At a stable equilibrium $\boldsymbol{\theta}$, the last block $U_{k}$ must synchronize. This follows directly from Lemma 5.2, as the nodes in $U_{k}$ are false twins and thus must synchronize.
Inductive hypothesis. Assume that the nodes in the last $m$ blocks of 1’s, $U_{k-m+1},\dots,U_{k}$, are synchronized.
Inductive step. Here, we show that the synchronization of $\bigcup_{i=k-m+1}^{k}U_{i}$ propagates to $\bigcup_{i=k-m}^{k}U_{i}$, since the preceding block $U_{k-m}$ must synchronize with $\bigcup_{i=k-m+1}^{k}U_{i}$. Indeed, since all nodes in $\bigcup_{i=k-m+1}^{k}U_{i}$ synchronize, any node added as isolated nodes whose neighbourhood is a subset of $\bigcup_{i=k-m+1}^{k}U_{i}$ must synchronize with them as well. These nodes are exactly nodes in $\bigcup_{i=k-m+1}^{k}I_{i}$. That is, every nodes $i,j\in\bigcup_{i=k-m+1}^{k}U_{i}\cup I_{i}$ satisfy $\boldsymbol{v}_{i}=\boldsymbol{v}_{j}$.
Observe that $\bigcup_{i=k-m+1}^{k}I_{i}$ are in fact synchronizing pendant attaches to $\bigcup_{i=k-m+1}^{k}U_{i}$, and they are the only difference of neighbourhood between the $U_{k-m}$ and $\bigcup_{i=k-m+1}^{k}U_{i}$, thus according to Lemma 6.6, nodes in $\bigcup_{i=k-m}^{k}U_{i}$ form a set of geometric stable closed twins, and they must synchronize.
Conclusion. All nodes in $\cup_{i=1}^{k}\left(U_{i}\cup I_{i}\right)$ or $I_{0}\cup\left(\cup_{i=1}^{k}(U_{i}\cup I_{i})\right)$ synchronize. Namely $\theta_{1}=\cdots=\theta_{n}$, thus any threshold graph is globally synchronizing as long as it is connected. ∎

### v2

@@ 5.3. Synchronous Pendant Extension
In this section, we present a scenario in which a group of nodes satisfy the phasor relations of Lemma 5.1 without being closed twins. More explicitly, the main result of this section is Lemma 5.8, which shows that even when a set of nodes are not twins, they may still be forced to synchronize at any second-order stationary point, provided that their distinct neighbors form a synchronous pendant whose combined phasor contribution collapses to a single, non-adversarial direction.
The next lemma, Lemma 5.7, serves as a preparatory step toward establishing Lemma 5.8.
@@
Let $G=(V,E)$ be a graph, and let $W\subseteq V$ be partitioned into
$W=Q\uplus S\uplus P,$
subject to the following structural conditions:
  1. (i)
    Every node in $Q$ has all its neighbors inside $S$, i.e., $N(i)\subseteq S$ for all $i\in Q$;
  2. (ii)
    $S$ induces a clique in $G$;
  3. (iii)
    For each $i\in S$, its neighborhood is $P\cup N_{Q}(i)\cup(S\setminus\{i\})$, where $N_{Q}(i)=\{\,q\in Q:q\sim i\,\}$.
Assume further that the configuration $(\boldsymbol{v}_{i})_{i\in V}$ is a second-order stationary point of the Kuramoto energy (4), and that $\boldsymbol{v}_{i}=\boldsymbol{v}$ for all $i\in Q\uplus S$. Then the phasor sum over $P$ satisfies
$\sum_{i\in P}\boldsymbol{v}_{i}\;\Uparrow\;\boldsymbol{v}\qquad\text{or}\qquad\sum_{i\in P}\boldsymbol{v}_{i}=\boldsymbol{0}.$
@@
$\sum_{j\in Q\uplus(S\setminus\{i\})\uplus P}\boldsymbol{v}_{j}=\mu_{i}\boldsymbol{v}_{i}.$
Since $Q\uplus S$ forms a synchronized group, we have
$\bigl(|Q|+|S|-1\bigr)\,\boldsymbol{v}_{i}+\sum_{j\in P}\boldsymbol{v}_{j}=\mu_{i}\boldsymbol{v}_{i},\qquad\mu_{i}\geq 0.$
@@
$\sum_{j\in P}\boldsymbol{v}_{j}=\bigl(\mu_{i}-(|Q|+|S|-1)\bigr)\boldsymbol{v}_{i},\qquad\mu_{i}\geq 0.$
We now show that $\mu_{i}-(|Q|+|S|-1)<0$ is impossible, because in this case there would exist a direction along which the energy decreases in a neighborhood of $\boldsymbol{\theta}$.
To see this, let $\boldsymbol{v}$ be the $n$-dimensional indicator vector of $Q\uplus S$, i.e., $v_{i}=1$ for $i\in Q\uplus S$ and $v_{i}=0$ otherwise. Then
$\boldsymbol{v}^{T}H\boldsymbol{v}$ $=\sum_{i\in Q\uplus S}\sum_{j\in P}\cos(\theta_{i}-\theta_{j})$ (17)
@@
$=|S|\,\boldsymbol{v}_{i}\cdot\sum_{j\in P}\boldsymbol{v}_{j}.$
where the second equality uses the fact that there are no edges between $Q$ and $P$. Clearly, if $\mu_{i}-(|Q|+|S|-1)<0$, then the above quantity is negative, which contradicts the second-order stationarity of $\boldsymbol{\theta}$. ∎
Figure 11. Illustration of Lemma 5.7. The green vertices correspond to nodes for which $\mu_{i}\boldsymbol{v}_{i}=\sum_{j\in N(i)}\boldsymbol{v}_{j}$ holds with $\mu_{i}\geq 0$. The large circular node represents the set $P$ of vertices. Gray edges indicate synchronization interactions between the corresponding nodes.
@@
(ii) The subgraph $G[S_{1}\uplus S_{2}]$ is a complete graph;
(iii) For any $i\in S_{1}\uplus S_{2}$, its neighbourhood can be decomposed into three parts: the set $P$, its neighbours inside $Q$, and the rest of the clique $S_{1}\uplus S_{2}$. Formally, if we denote $Q_{i}=N(i)\cap Q$, then
$N(i)\;=\;P\uplus Q_{i}\uplus\bigl((S_{1}\uplus S_{2})\setminus\{i\}\bigr).$
@@
$N(i)\setminus\bigl(Q\uplus(S_{1}\uplus S_{2})\bigr)\;=\;P.$
Assume further that all nodes in $Q\uplus S_{1}$ are synchronized, i.e.,
$\boldsymbol{v}_{i}=\boldsymbol{v}_{j},\quad\forall i,j\in Q\uplus S_{1}.$
Then the set $S_{1}\uplus S_{2}$ synchronize.
@@
$\forall i\in S_{1},\quad\boldsymbol{v}_{i}\Uparrow\sum_{k\in S_{2}\uplus P}\boldsymbol{v}_{k}\quad\text{or}\quad\sum_{k\in S_{2}\uplus P}\boldsymbol{v}_{k}=\boldsymbol{0}.$
@@
$\forall i\in S_{1},\quad\boldsymbol{v}_{i}\Uparrow\sum_{k\in S_{2}\uplus P}\boldsymbol{v}_{j}+\sum_{k\in S_{1}\setminus\{i\}}\boldsymbol{v}_{j}\quad\text{or}\quad\sum_{k\in S_{2}\uplus P}\boldsymbol{v}_{k}+\sum_{k\in S_{1}\setminus\{i\}}\boldsymbol{v}_{k}=\boldsymbol{0}.$
@@
$\forall i\in S_{1},\quad\boldsymbol{v}_{i}\Uparrow\sum_{k\in S_{1}\uplus P\uplus S_{2}\setminus\{i\}}\boldsymbol{v}_{k}\quad\text{or}\quad\sum_{k\in S_{1}\uplus P\uplus S_{2}\setminus\{i\}}\boldsymbol{v}_{k}=\boldsymbol{0}.$ (18)
@@
$\forall i\in S_{2},\quad\boldsymbol{v}_{i}\Uparrow\sum_{k\in S_{2}\setminus\{i\}}\boldsymbol{v}_{k}+\sum_{j\in S_{1}\uplus P}\boldsymbol{v}_{k}\quad\text{or}\quad\sum_{k\in S_{2}\setminus\{i\}}\boldsymbol{v}_{k}+\sum_{j\in S_{1}\uplus P}\boldsymbol{v}_{k}=\boldsymbol{0}$
@@
$\forall i\in S_{2},\quad\boldsymbol{v}_{i}\Uparrow\sum_{k\in S_{1}\uplus P\uplus S_{2}\setminus\{i\}}\boldsymbol{v}_{k}\quad\text{or}\quad\sum_{k\in S_{1}\uplus P\uplus S_{2}\setminus\{i\}}\boldsymbol{v}_{k}=\boldsymbol{0}.$ (19)
Combined the facts (18) and (19), we get nodes in $S_{1}\uplus S_{2}$ are geometric stable closed twins. ∎
Clearly, this is another situation in which geometric twins, as defined in Definition 5.3, appear in addition to closed twins at the second-order stationary point.
Figure 12. Illustration of Lemma 5.8, illustrating how the synchronized block $Q\uplus S_{1}$ extends to include $S_{2}$ through propagation enabled by the specific local structure. The green vertices correspond to nodes for which $\mu_{i}\boldsymbol{v}_{i}=\sum_{j\in N(i)}\boldsymbol{v}_{j}$ holds with $\mu_{i}\geq 0$. The gray-shaded nodes are synchronized; an edge with a gray background indicates synchronization between the corresponding nodes. Each orange arc separates a small green node from a group of nodes it connects to, indicating that the vector of the green node is in the same direction as the sum of the vectors of the nodes on the opposite side of the arc.
5.4. Twin Attachment
If a set of nodes all have identical neighborhoods, they form a group of closed twins. When an additional node is connected exclusively to this group, the entire collection of nodes, must take the same phase at any second-order stationary point.
Lemma 5.9.
Let $\boldsymbol{\theta}$ be a second-order stationary point of $E_{G}(\boldsymbol{\theta})$. If $S\subseteq V$ is a set of closed twins in $G$ and $A\in V$ satisfies $N(A)=S$, then $\theta_{i}=\theta_{j}$ for all $i,j\in S\cup\{A\}$.
Proof.
Since $\boldsymbol{\theta}$ is a second-order stationary point of $E_{G}(\boldsymbol{\theta})$ and $S$ forms a set of closed twins, Lemma 5.5 yields
$\theta_{i}=\theta_{j}\qquad\text{for all }i,j\in S.$
Because $N(A)=S$, Lemma 4.4 implies
$\theta_{A}=\theta_{i}\qquad\text{for all }i\in S.$
Hence all nodes in $S\cup\{A\}$ synchronize. ∎
@@ 6. Main Proof: the Induction Argument
6.1. A Prototype Example
Let us take a look at the threshold graph corresponding to the sequence 01010101. The first vertex is denoted by $A$, and we label the vertices $A$ through $I$ in the order in which they appear in the sequence, with each bit determining whether the added vertex is isolated (0) or dominating (1).
Let $G$ be the threshold graph corresponding to the sequence 01010101. To prove that $G$ is globally synchronizing, we show that every second-order stationary point $\boldsymbol{\theta}$ of $E_{G}(\boldsymbol{\theta})$ is synchronous; that is, $\theta_{1}=\cdots=\theta_{n}$. The following figure illustrates how synchronization must propagate, step by step, in any second-order stationary configuration on this graph.
ACEGIHFDB
ACEGIHFDB
ACEGIHFDB
ACEGIHFDB
ACEGIHFDB
ACEGIHFDB
ACEGIHFDB
ACEGIHFDB
Figure 14. Synchronization propagates in eight steps, forcing every second-order stationary point to be synchronous.
Step (1) We begin with the edge between $H$ and $I$. Applying Lemma 4.4 to node $H$ shows that, since $\boldsymbol{\theta}$ is a second-order stationary point, we must have $\boldsymbol{v}_{H}\Uparrow\boldsymbol{v}_{I}$. Equivalently, nodes $H$ and $I$ synchronize.
Step (2) Observe that $N(G)=N(I)\cup H$ and $\theta_{H}=\theta_{I}$, applying Lemma by letting $Q=\{H\}$, $S_{1}=\{I\}$, $S_{2}=\{G\}$, $P=\{F,D,B,A,C,E\}$, we obtain $\theta_{G}=\theta_{I}$. Intuitively, the only difference between the neighborhoods of $G$ and $H$ is that $H$ has one additional neighbor, namely $I$. This extra neighbor does not obstruct the synchronization between $G$ and $I$, because $H$ is already synchronized with $I$ and therefore behaves as a synchronous pendant at this stage.
@@ 6.1. A Prototype Example
Step (3) Note that node $F$ connects only to $G$ and $I$, which are synchronized as established in Step (2). Therefore, by Lemma 5.9, we have $\boldsymbol{v}_{F}=\boldsymbol{v}_{G}=\boldsymbol{v}_{I}$.
Step (4) Observe that nodes $F$ and $H$ connect only to $G$ and $I$, and the four-node set $\{F,H,G,I\}$ behaves as a synchronizing group. Furthermore, ignoring the leaf-like nodes $F$ and $H$ and edges in the induced clique on $\{E,G,I\}$, the two nodes $G$ and $I$ share exactly the same neighbors:
$N(G)\setminus(\{F\}\cup\{E,I\})\;=\;N(I)\setminus(\{F,H\}\cup\{E,G\})\;=\;N(E)\setminus\{G,I\}\;=\;\{D,B,A,C\}.$
Therefore, by applying Lemma 5.8 with $Q=\{F,H\}$, $S_{1}=\{G,I\}$, $S_{2}=\{E\}$, and $P=\{D,B,A,C\}$, we conclude that the nodes $E$, $G$, and $I$ form a synchronizing group.
@@
Step (5) Note that node $D$ connects to $E$, $G$, and $I$, which are synchronized as established in Step (4). Therefore, by Lemma 5.9, we have $\boldsymbol{v}_{D}=\boldsymbol{v}_{E}=\boldsymbol{v}_{G}=\boldsymbol{v}_{I}$.
Step (6) Observe that nodes $D$, $F$, and $H$ connect only to $E$, $G$, and $I$, and therefore the six-node set $\{D,F,H,E,G,I\}$ behaves as a synchronizing group. Furthermore, ignoring the leaf-like nodes $D$, $F$, and $H$ and edges in the induced clique on $\{C,E,G,I\}$, the four nodes $C$, $E$, $G$, and $I$ share exactly the same neighbors, namely $\{A,B\}$. Therefore, by applying Lemma 5.8 with $Q=\{D,F,H\}$, $S_{1}=\{E,G,I\}$, $S_{2}=\{C\}$, and $P=\{A,B\}$, we conclude that the nodes $C$, $E$, $G$, and $I$ form a synchronizing group.
@@
Step (7) Note that node $B$ connects to $C$, $E$, $G$, and $I$, which are synchronized as established in Step (6). Therefore, by Lemma 5.9, we have $\boldsymbol{v}_{B}=\boldsymbol{v}_{C}=\boldsymbol{v}_{E}=\boldsymbol{v}_{G}=\boldsymbol{v}_{I}$.
Step (8) As in Step (7), Lemma 5.9 implies that $A$ synchronizes with $C$, $E$, $G$, and $I$, since these nodes form a synchronizing group and constitute all neighbors of $A$.
In conclusion, $\boldsymbol{\theta}$ must assign the same angle to every node. Equivalently, every second-order stationary point is synchronous, and hence the threshold graph is globally synchronizing.
6.2. Proof of Theorem 1.2: General Case
Proof of Theorem 1.2.
@@
Any connected threshold graph can be described by a bit sequence of $1$’s and 0’s, as introduced in Definition 2.1. Since the graph is connected, the last bit must be a $1$. We decompose the sequence into maximal contiguous blocks of $1$’s and 0’s. Let $U_{1},U_{2},\dots,U_{k}$ denote the blocks of $1$’s (indexed from bottom to top), and let $I_{i}$ denote the 0-block that lies immediately above $U_{i}$, if such a block exists. In this notation, the sequence has one of the two forms
@@
depending on whether or not it begins with a block of 0’s. We focus on the first case; the second follows by an entirely analogous argument.
Given that $\boldsymbol{\theta}$ is a second-order stationary point of $E(\boldsymbol{\theta})$, we aim to show that $\theta_{1}=\cdots=\theta_{n}$, and we proceed by induction. We proceed by induction. For the base case, observe that all nodes within a block of $1$’s form a set of closed twins. Hence, at any second-order stationary point $\boldsymbol{\theta}$, the nodes in the last block $U_{k}$ must synchronize, as follows directly from Lemma 5.1.
Assume as the inductive hypothesis that the nodes in the last $m$ blocks of $1$’s, $U_{k-m+1},\dots,U_{k}$, are synchronized. Lemma 5.9 then implies that the adjacent 0-blocks $I_{k-m+1},\dots,I_{k}$ synchronize with them as well. We next show that the preceding $1$-block $U_{k-m}$ also synchronizes with $U_{k-m+1},\dots,U_{k}$. Indeed, define
$Q$ $=I_{k-m+1}\uplus\cdots\uplus I_{k},$
$S_{1}$ $=U_{k-m+1}\uplus\cdots\uplus U_{k},$
$S_{2}$ $=U_{k-m},$
$P$ $=U_{1}\uplus\cdots\uplus U_{k-m-1}\uplus I_{2}\uplus\cdots\uplus I_{k-m}.$
We next verify that the sets $Q,S_{1},S_{2}$, and $P$ satisfy the assumptions of Lemma 5.8. First, for every node $i$ in $Q$, its neighborhood is contained in $S_{1}$; indeed, these nodes are isolated at this stage and connect only to the dominating nodes that are added later. Second, the sets $S_{1}$ and $S_{2}$ together induce a clique. Third, for any node in $S_{1}\cup S_{2}$, its neighbors consist exactly of all other nodes in $S_{1}\cup S_{2}$, all nodes in $P$, and possibly some nodes in $Q$. Hence all conditions of Lemma 5.8 are satisfied, and we conclude that $S_{1}$ and $S_{2}$ must synchronize. In particular, the block $U_{k-m}$ synchronizes with $U_{k-m+1}\uplus U_{k-m+2}\uplus\cdots\uplus U_{k}$.
By induction, all nodes synchronize at $\boldsymbol{\theta}$. Therefore, every second-order stationary point is a synchronous state. It follows that every connected threshold graph is globally synchronizing. ∎
7. Discussion and Open Problems
In this work, we present a class of globally synchronizing graphs — threshold graphs — whose density vary from $2/n$ to $1$, revealing a new mechanism of global synchronization that has not been articulated before. Our proof crucially exploits the constructive nature of threshold graphs: they form the smallest graph class closed under the addition of isolated nodes and universal nodes. This closure property is essential to our inductive argument, as it allows us to show that global synchronization is preserved when starting from a threshold graph and adding either an isolated node or a universal node.
A natural question is whether this closure phenomenon extends beyond threshold graphs. Specifically, suppose $G$ is an arbitrary globally synchronizing graph. If we add a universal node to $G$, is the resulting graph still globally synchronizing? We expect the answer to be yes, since attaching a node that is connected to all existing vertices does not appear to introduce new geometric directions for the oscillators beyond the synchronized configuration. However, we do not currently know how to prove this, and we leave this problem for future work.
Moreover, we expect that our results may admit further generalizations. Possible directions include extending the analysis to oscillators evolving on higher-dimensional spheres, incorporating nonlinear interaction terms as in [Geshkovski2025], and considering broader classes of graphs. For instance, one may study systems in which oscillators are coupled through both positive and negative edges.